OpenStudy (anonymous):

(2,1.5),(3,4) Find slope

OpenStudy (anonymous):

2.5

OpenStudy (anonymous):

Slope is rise over run. The difference in y values divided by the difference in x values: (4 - 1.5)/(3 - 2) = 2.5/1 = 2.5

OpenStudy (jamesj):

http://www.youtube.com/watch?v=_4hUYnLGN7g